OpenStudy (anonymous):

Brandon can paint a fence in 12 hours and Elaine can paint fence in 11 hours.How long will they take to paint the fence if they work together

OpenStudy (anonymous):

\[\frac{11\times 12}{11+12}\] is the short quick answer

OpenStudy (anonymous):

we can work through all the steps, but in the end that is what you will get

OpenStudy (anonymous):

how

OpenStudy (anonymous):

brandon's rate is \(\frac{1}{12}\) of the fence per hour elaine's rate is \(\frac{1}{11}\) of the fence per hour their combined rate is the sum of these two, namely \[\frac{11+12}{11\times 12}\] you want to see how long it takes to complete one job, so set \[\frac{11+12}{11\times 12}t=1\] and solve for \(t\) in one step: \[t=\frac{11\times 12}{11+12}\] as promised

OpenStudy (anonymous):

at the first step when you add \[\frac{1}{12}+\frac{1}{11}\] you can go ahead and write \[\frac{1}{12}+\frac{1}{11}=\frac{23}{132}\] then when you solve \[\frac{23}{132}t=1\] you get \[t=\frac{132}{23}\] but if you do more than one of these, or if you have a test, you can do it quickly as i wrote above

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

yw